5.Small aircrafts arrive at a certain airport at a rate of 5 per hour. What is the probability that an aircraft arrive in the next 2 hours?
Answer #1

Ans:

Use exponential distribution:

lambda=5 per hr

P(T<=2)=1-e-(5*2)=1-e-10=0.999955